# ZeroFlow: Overcoming Catastrophic Forgetting is Easier than You Think

### v2

@@ Abstract
Backpropagation provides a generalized configuration for overcoming catastrophic forgetting. Like, SGD and Adam are commonly used for weight updates in continual learning and continual pre-training. In practice, permission to access gradient information is not always granted (the gradient ban), such as black-box APIs, hardware limitations, and non-differentiable systems. To bridge this gap, we introduce the first benchmark ZeroFlow to evaluate gradient-free optimization algorithms for overcoming forgetting. This benchmark examines a suite of forward pass methods across multiple methods, forgetting scenarios, and datasets. We find that forward passes alone are enough to overcome forgetting. Our findings reveal new optimization principles that highlight the potential of forward-pass in mitigating forgetting, managing task conflicts, and reducing memory demands, alongside novel enhancements that further mitigate forgetting with just one forward pass. This work provides essential insights and tools for advancing forward pass methods to overcome forgetting.
@@ 1 Introduction
Catastrophic forgetting remains one of the major challenges on the path to artificial general intelligence (AGI) [21, 58, 40, 67], i.e., models tend to forget previously learned tasks when trained on new ones on time-evolving data flow [21, 8, 13]. This phenomenon is commonly seen across various tasks [13], including continual learning (CL) [67, 58], fine-tuning of foundation models (FMs) [39, 71, 63], and continual pre-training (CPT) [37, 49, 72], etc. Among them, optimization algorithms play a crucial role, e.g., SGD has become the default choice during CL [57, 51], while Adam is frequently seen in fine-tuning foundation models [37, 71]. These optimization algorithms in tandem with various methods (ranging from regularization and rehearsal strategies to architectural changes) rely on gradient information to avoid forgetting [68, 3, 36, 24]. Nonetheless, in real-world scenarios, gradient information is not always available or computable (i.e., the gradient ban), like,
Figure 1: Illustrations of ZeroFlow. New tasks (or downstream tasks) arrive sequentially, the gradient bans block the model from learning and memorizing using backpropagation. Our ZeroFlow overcome this issue via forward passes.
Scenario i: large language models as a service (LLMaaS) and black-box APIs.
Scenario ii: hardware systems that do not support principled backpropagation.
Scenario iii: AI for science with non-differentiable underlying systems.
In other words, Scenario i implies that pretrained models are monetized [10, 52, 42] (model owners do not publicly release their pretrained models but instead the service), i.e., only the inputs and outputs are accessible [18, 53]. Scenarios ii/iii implies that the limitations prevent or restrict the execution of backpropagation [32, 59], i.e., extremely high memory demands [39, 24], unsupported systems and hardware [25], or non-differentiable functions, etc [55, 20]. The above means that typical methods for overcoming forgetting are not available because backpropagation is banned, as Figure 1. This yields the primary question to be explored,
@@
Figure 2: ZeroFlow Evaluation Results of Catastrophic Forgetting. We visualize the evaluation results of 2 models in several ZeroFlow dimensions (average accuracy over all tasks and a forgetting metric). For comprehensive numerical results, please refer to Table 1.
To tackle (Q), a natural idea is to use the forward pass method [23, 1, 44] instead of backpropagation to overcome forgetting. Such as the zeroth-order (ZO) optimization method [15, 43, 38, 19], category of methods is well-suited to this issue due to their relaxed information requirements (requiring only the function values rather than gradients). Under gradient bans, DECL and DFCL first attempt to overcome forgetting from a stream of APIs [61], but they focus on synthetic data level rather than optimization. Therefore, it remains elusive whether benchmark studies using gradient-free methods can mitigate catastrophic forgetting.
@@
(iii) Apart from a comprehensive evaluation of catastrophic forgetting, we introduce two enhancement techniques, which further improve the performance and efficiency of just one forward pass to overcome forgetting.
@@ 2 Literatures
2.1 Catastrophic forgetting
Catastrophic forgetting occurs across various tasks, including CL, fine-tuning of FMs, and CPT [67, 58, 71, 37]. To mitigate this issue, various methods have been proposed [45, 26, 54, 30]. In CL, methods range from regularization and rehearsal strategies to architectural changes [28, 3, 36]. Lately, pre-trained models (PTM) further advanced these methods due to their strong generalization [62, 12], as seen in PTM-based CL [69]. All these methods share a common goal: achieving an optimal balance between learning plasticity and memory stability [58]. In FMs, catastrophic forgetting often arises from overfitting to small fine-tuning datasets during CPT or fine-tuning [37, 71]. Common techniques to address this include learning rate adjustment, parameter-efficient fine-tuning, mixed data strategies, and instruction tuning [37, 24]. Additionally, as foundational models increasingly gain multimodal capabilities, the complexity of catastrophic forgetting also intensifies [71].
2.2 Optimization for catastrophic forgetting
Two broad categories of optimization methods exist for overcoming forgetting, (i) Standard Optimization. SGD and the Adam family are frequently employed to investigate catastrophic forgetting [4, 21, 40]. For instance, in CL, various CL methods predominantly utilize the SGD optimizer for standard evaluations [57, 51, 70]. In fine-tuning the LLM, the Adam series is commonly used to observe forgetting phenomena [37, 71]. Some works explored orthogonal spaces with these standard optimizers to alleviate forgetting [35, 14, 47], such as OGD [11], and GPM [47]. Moreover, other works [11, 5, 35] modified the gradients in the standard optimization process to align the learning spaces of new and old tasks, such as Uni-Grad [30]. The core of these efforts [9, 48] is to find an equilibrium between learning and forgetting in optimization. (ii) Sharpness-aware Optimization. This series of methods [22, 16, 65, 73] started to draw attention due to the effectiveness of the flat minimum in mitigating forgetting [30, 29, 4, 41]. For instance, FS-DPGM [9], F2M [48], DFGP [60], SAM-CL [56], etc. overcome forgetting in the flatness areas of different configurations. C-Flat [3] proposed a CL-friendly general optimization framework, that holds promise as a baseline optimizer for overcoming catastrophic forgetting.
@@ 3.1 Zeroth-order Optimization
1 Input Initialized model parameters $\mathbf{\theta}_{0}\in\Theta\subseteq\mathbb{R}^{d}$, training dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{m}\in\mathcal{X}\times\mathcal{Y}$, empirical loss function $\mathcal{L}$, learning rate $\eta_{t}$, gradient perturbation vector $\mathbf{\xi}$, and descent direction computation $\phi(\cdot)$
2 while $\theta_{t}$ not converged do
3       Sample mini-batch $\mathcal{B}$ from $\mathcal{D}$
4       Step 1. $\triangleright$ ZO gradient estimation:
5       $\hat{\mathbf{g}}_{t}=\hat{\nabla}\mathcal{L}(\theta,\mathbf{\xi};\mathcal{B})$
6       Step 2. $\triangleright$ Descent direction computation:
7       $\mathbf{h}_{t}=\phi\left(\{\hat{\mathbf{g}}_{i}\}_{i=1}^{t}\right)$
8       Step 3. $\triangleright$ Parameter updating:
9       $\mathbf{\theta}_{t+1}=\mathbf{\theta}_{t}-\eta_{t}\cdot\mathbf{h}_{t}$
10       $t=t+1$
Output Updated model $\mathbf{\theta}_{t}$
@@ 10       $t=t+1$
1) ZO gradient estimation. Randomized Gradient Estimation (RGE [43]) and Coordinate-wise Gradient Estimation (CGE [2]) perturb the model using $\mathcal{\xi}$, which is generated either from a random unknown distribution (in RGE) or by modifying individual coordinates (in CGE), and then observe the changes in the loss function $\mathcal{L}$ after each perturbation, step by step, to provide a reliable gradient estimate. However, due to their reliance on slow single-direction perturbation, these methods are not well-suited for deep learning tasks, as performing a full perturbation in high-dimensional parameter spaces is time-consuming. For instance, the ImageNet dataset has input dimensions of 224×224×3, totaling approximately 0.13 million parameters, and rendering over each dimension iteratively leads to query inefficiency. The standard Simultaneous Perturbation Stochastic Approximation (SPSA) improves efficiency by generating pairs of symmetric vectors and perturbing in multiple directions simultaneously, as follows
@@
Where $\epsilon$ is a positive scaler and $\xi$ is recommended to follow a symmetric distribution with finite inverse moments (e.g., the Rademacher distribution). The symmetric distribution ensures unbiased exploration of perturbations in both positive and negative directions of parameters at each step. And the finite inverse moments property guarantees that the steps are well-controlled, avoiding excessively large steps due to $\xi^{-1}$ drawn from the distribution (e.g., $\mathbb{E}[1/|\xi|^{p}]$ for some large p), which would otherwise lead to an unstable optimization process. In practical implementations for models with a large number of parameters (e.g., MeZO [38] in large language models (LLMs)), Gaussian noise with zero mean induces substantial perturbations, thereby enhancing exploration across the parameter space and facilitating the escape from local minima. This methodology achieves gradient estimation with only two objective function evaluations, rendering its computational cost independent of input dimensionality. Such computational efficiency has established SPSA as a preferred method for addressing the complexities of high-dimensional deep learning tasks. While increasing $q$ in $q$-SPSA can improve stability in the update direction, setting $q=1$ is sufficient for pretrained LLMs  [38].
2) Descent direction computation. In unconstrained optimization for deep learning, the last gradients $h_{t}$ generally coincide with the estimated ZO gradients $\hat{g_{t}}$ (e.g., ZO-SGD [19], ZO-SCD [31]). To reduce approximation errors, ZO-SGD-sign [34] applies an element-wise $sign(\cdot)$ operation. Additionally, ZO-SVRG [33], inspired by variance reduction methods in first-order optimization, adjusts the update step by using estimated gradients from previous training examples. The CARS method [27] adaptively selects the smallest function value in each iteration, which helps maintain monotonicity in the optimization process.
3) Parameter updating. Normally, for most ZO methods, parameters are updated in a similar way with FO optimizers, and the learning rate $\eta_{t}$ is set to constant. Except for the special design for achieving some constraint prerequisites, several methods make an effort to strike a balance between converge speed and accuracy. ZO-AdaMM [7] uses an adaptive learning rate and refines gradient estimation by incorporating momentum from past information. This approach is particularly effective in handling complex and evolving optimization landscapes, where the function’s behavior may vary over time or be hard to capture with straightforward gradient approximations.
@@ 3.2 Zeroth-order Optimization for Catastrophic Forgetting
Rationality. ZO optimization leverages the function values of the forward passes to approximate FO gradients, making it feasible to avoid gradient bans. This feature enables seamless integration into common forgetting scenarios, such as continual learning. We explore it in the following three categories.
@@
where $N_{context}$ represents the number of contexts encountered so far. In Experience Replay [46], both components use classification loss based on their respective data distributions, so ZO gradients can be expressed as ^$\hat{\nabla}\mathcal{L}_{cur}$ and ^$\hat{\nabla}\mathcal{L}_{replay}$ respectively. However, in the emerging generative replay workflows [50], Eq. 2 may introduce additional loss for the training of generators. In this case, the generator can be trained using standard backpropagation or in conjunction with ZO training without FO gradients.
ii) Extension-based methods can be divided into fixed and dynamic architectures. Fixed architectures separate model parameters for specialized context learning, while dynamic architectures expand the model size during adaptation. Both approaches mitigate catastrophic forgetting from the model’s perspective and enable model-agnostic ZO solutions.
@@
(a) FO-SGD
(b) ZO-SGD
Figure 3: The Trajectory of FO and ZO Optimization during Overcoming Forgetting. The trajectory has taken when using the total loss from both tasks (cyan) and the gradients from each individual task at fixed points during optimization (red and blue). The trends of ZO optimization hold the potential to manage forgetting and learning.
As shown in Figure 3, we visualize and compare the optimization trajectories of ZO and FO methods in the learning-memory trade-off dynamics of continual learning. The remarkable similarity between these trajectories motivates our investigation, suggesting ZO optimization’s potential in effectively managing the balance between learning and forgetting.
Potential. The intrinsic optimization mechanism of ZO exhibits particular promise in continual learning scenarios. Intuitively, ZO perturbs parameters using random or coordinate-wise directional vectors and observes changes in the evaluation function, effectively optimizing within a noisy environment. This approach enables small parameter modifications to yield significant impacts on target objectives, resulting in distinctive gradient estimations compared to FO optimization. Notably, while ZO methods do not explicitly incorporate sharpness regularization terms [17], they naturally facilitate the exploration of flat regions in parameter space. The influence of optimizing flat regions with ZO approaches in continual learning can be summarized in two main manifolds: (i) For previous tasks, the noise-induced parameter robustness enhances resilience against perturbations from new task adaptation; (ii) For new tasks, empirical evidence suggests that convergence to flat minima generally leads to lower generalization error.
Table 1: ZeroFlow Evaluation Results on CIFAR-100, ImageNet-A, CUB and OmniBenchmark. This table compares the average accuracy, final accuracy, and forgetting measures of two models, and four forgetting scenarios. More intuitive trend please see Figure 4. Note that all ZO optimization w/ the query budget $q=1$, which means that all our evaluations use just one forward pass.
@@
Risk. Although ZO demonstrates superior generalization abilities, its practical performance is limited by optimization strategies and the complexity of the optimization setting. Despite significant efforts to reduce convergence error, optimizing models from scratch in high-dimensional space remains challenging due to slow convergence speed (proportional to the parameter dimension $d$). For instance, origin CGE-based ZO training for a model with 12k parameters takes 70.32 hours in DeepZero [6]. Such computational demands render from-scratch training impractical for high-dimensional continual learning models, particularly those employing expansion-based architectures. Consequently, we focus our discussion on leveraging ZO optimization to overcome forgetting within a pre-training context.

### v3

@@ Abstract
Backpropagation provides a generalized configuration for overcoming catastrophic forgetting. Like, SGD and Adam are commonly used for weight updates in continual learning and continual pre-training. In practice, permission to access gradient information is not always granted (the gradient ban), such as black-box APIs, hardware limitations, and non-differentiable systems. To bridge this gap, we introduce the first benchmark ZeroFlow to evaluate gradient-free optimization algorithms for overcoming forgetting. This benchmark examines a suite of forward pass methods across multiple methods, forgetting scenarios, and datasets. We find that forward passes alone are enough to overcome forgetting. Our findings reveal new optimization principles that highlight the potential of forward-pass in mitigating forgetting, managing task conflicts, and reducing memory demands, alongside novel enhancements that further mitigate forgetting with just one forward pass. This work provides essential insights and tools for advancing forward pass methods to overcome forgetting. Code will be available upon publication.
Machine Learning, ICML
@@ 1 Introduction
Catastrophic forgetting remains one of the major challenges on the path to artificial general intelligence (AGI) (Hadsell et al., 2020; Zhou et al., 2023b), i.e., models tend to forget previously learned tasks when trained on new ones on time-evolving data flow (Feng et al., 2022b). This phenomenon is commonly seen across various tasks, including continual learning (CL) (Wang et al., 2023), fine-tuning of foundation models (FMs) (Mangrulkar et al., 2022; Yuan et al., 2024), and continual pre-training (CPT) (Shi et al., 2024; Zhu et al., 2024b), etc. Among them, optimization algorithms play a crucial role, e.g., SGD has become the default choice during CL (van de Ven et al., 2022), while Adam is frequently seen in fine-tuning FMs (Luo et al., 2023; Zhu et al., 2024a). These optimization algorithms in tandem with various methods (ranging from regularization and rehearsal strategies to architectural changes) rely on gradient information to avoid forgetting (Zhou et al., 2023c; Bian et al., 2024). Nonetheless, in real-world scenarios, gradient information is not always available or computable (i.e., the gradient ban), like, Scenario i: large language models as a service (LLMaaS) and black-box APIs. Scenario ii: hardware systems that do not support principled backpropagation. Scenario iii: AI for science with non-differentiable underlying systems.
Figure 1: Illustrations of ZeroFlow. New tasks (or downstream tasks) arrive sequentially, the gradient bans block the model from learning and memorizing using backpropagation. ZeroFlow overcome this issue via forward passes.
In other words, Scenario i implies that pretrained models are monetized (Miura et al., 2024) (model owners do not publicly release their pretrained models but instead the service), i.e., only the inputs and outputs are accessible (Gan et al., 2023; Sun et al., 2022). Scenarios ii/iii implies that the limitations prevent or restrict the execution of backpropagation (Lillicrap et al., 2020), i.e., extremely high memory demands (Mangrulkar et al., 2022), unsupported systems and hardware (Jabri & Flower, 1992), or non-differentiable functions, etc (Tavanaei et al., 2019; Gu et al., 2021). The above means that typical methods for overcoming forgetting are not available because backpropagation is banned, as Figure 1. This yields the primary question to be explored,
@@
Figure 2: ZeroFlow Evaluation Results of Catastrophic Forgetting. We visualize the evaluation results of 2 models (EASE (Zhou et al., 2024b) and APER (Zhou et al., 2023a)) in several ZeroFlow dimensions (average accuracy over all tasks and a forgetting metric). For comprehensive numerical results, please refer to Table 1.
To tackle (Q), a natural idea is to use the forward pass method (Hinton, 2022; Baydin et al., 2022; Ren et al., 2022) instead of backpropagation to overcome forgetting. The zeroth-order (ZO) optimization methods (Flaxman et al., 2004; Nesterov & Spokoiny, 2017; Malladi et al., 2023; Ghadimi & Lan, 2013), as representative methods, are well-suited to this issue due to their relaxed information requirements, as they rely only on function values rather than gradients. Under gradient bans, DECL and DFCL (Yang et al., 2024) first attempt to overcome forgetting from a stream of APIs, but they focus on synthetic data level rather than optimization. Therefore, it remains elusive whether benchmark studies using gradient-free methods can mitigate forgetting.
@@
(iii) Apart from a comprehensive evaluation of catastrophic forgetting, we introduce three enhancement techniques, which further improve the performance and efficiency of just one forward pass to overcome forgetting.
@@ 2 Literatures
Catastrophic forgetting. Catastrophic forgetting occurs across various tasks, including CL, fine-tuning of FMs, and CPT (Zhou et al., 2023b; Wang et al., 2023; Zhu et al., 2024a; Luo et al., 2023). To mitigate this issue, various methods have been proposed (Rolnick et al., 2019a; Jeeveswaran et al., 2023; Sun et al., 2023b; Li et al., 2024). In CL, methods range from regularization and rehearsal strategies to architectural changes (Kirkpatrick et al., 2017; Bian et al., 2024; Lu et al., 2024). Lately, pre-trained models (PTM) further advanced these methods due to their strong generalization (Yuan et al., 2022; Feng et al., 2022a), as seen in PTM-based CL (Zhou et al., 2024a). All these methods share a common goal: achieving an optimal balance between learning plasticity and memory stability (Wang et al., 2023). In FMs, catastrophic forgetting often arises from overfitting to small fine-tuning datasets during CPT or fine-tuning (Luo et al., 2023; Zhu et al., 2024a). Common techniques to address this include learning rate adjustment, parameter-efficient fine-tuning, mixed data strategies, and instruction tuning (Luo et al., 2023; Hu et al., 2021). Additionally, as foundational models increasingly gain multimodal capabilities, the complexity of catastrophic forgetting also intensifies (Zhu et al., 2024a).
Optimization for catastrophic forgetting. Two broad categories of optimization methods exist for overcoming forgetting, (i) Standard Optimization. SGD and the Adam family are frequently employed to investigate catastrophic forgetting (Cha et al., 2021; Hadsell et al., 2020; Masana et al., 2022). For instance, in CL, various CL methods predominantly utilize the SGD optimizer for standard evaluations (van de Ven et al., 2022; Sun et al., 2023a; Zhou et al., 2024c). In fine-tuning the LLM, the Adam series is commonly used to observe forgetting phenomena (Luo et al., 2023; Zhu et al., 2024a). Some works explored orthogonal spaces with these standard optimizers to alleviate forgetting (Lopez-Paz & Ranzato, 2017; Feng et al., 2022c; Saha et al., 2020), such as OGD (Farajtabar et al., 2020), and GPM (Saha et al., 2020). Moreover, other works (Farajtabar et al., 2020; Chaudhry et al., 2018; Lopez-Paz & Ranzato, 2017) modified the gradients in the standard optimization process to align the learning spaces of new and old tasks, such as Uni-Grad (Li et al., 2024). The core of these efforts (Deng et al., 2021; Shi et al., 2021) is to find an equilibrium between learning and forgetting in optimization. (ii) Sharpness-aware Optimization. This series of methods (He et al., 2019; Foret et al., 2020; Zhong et al., 2022; Zhuang et al., 2022) has gained attention due to the effectiveness of the flat minimum in mitigating forgetting (Li et al., 2024; Kong et al., 2023; Cha et al., 2021; Mehta et al., 2023). Methods such as FS-DPGM (Deng et al., 2021), F2M (Shi et al., 2021), DFGP (Yang et al., 2023), SAM-CL (Tung et al., 2023) overcome forgetting in the flatness areas of different configurations. C-Flat (Bian et al., 2024) proposed a CL-friendly general optimization framework, that holds promise as a baseline optimizer for overcoming forgetting.
@@ 3.1 Zeroth-Order Optimization
0:  Initialized model parameters $\mathbf{\theta}_{0}\in\Theta\subseteq\mathbb{R}^{d}$, training dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{m}\in\mathcal{X}\times\mathcal{Y}$, empirical loss function $\mathcal{L}$, learning rate $\eta_{t}$, gradient perturbation vector $\mathbf{\xi}$, and descent direction computation $\phi(\cdot)$
1:  while $\theta_{t}$ not converged do
2:     Sample mini-batch $\mathcal{B}$ from $\mathcal{D}$
3:     Step 1. ZO gradient estimation:
4:     $\hat{\mathbf{g}}_{t}=\hat{\nabla}\mathcal{L}(\theta,\mathbf{\xi};\mathcal{B})$
5:     Step 2. Descent direction computation:
6:     $\mathbf{h}_{t}=\phi\left(\{\hat{\mathbf{g}}_{i}\}_{i=1}^{t}\right)$
7:     Step 3. Parameter updating:
8:     $\mathbf{\theta}_{t+1}=\mathbf{\theta}_{t}-\eta_{t}\cdot\mathbf{h}_{t}$
9:     $t=t+1$
10:  end while
10:  Updated model $\mathbf{\theta}_{t}$
1) ZO gradient estimation. Randomized Gradient Estimation (RGE (Nesterov & Spokoiny, 2017)) and Coordinate-wise Gradient Estimation (CGE (Berahas et al., 2022)) perturb the model using $\mathcal{\xi}$, which is generated either from a random unknown distribution (in RGE) or by modifying individual coordinates (in CGE), and then observe the changes in the loss function $\mathcal{L}$ after each perturbation, step by step, to provide a reliable gradient estimate. However, due to their reliance on slow single-direction perturbation, these methods are not well-suited for deep learning tasks, as performing a full perturbation in high-dimensional parameter spaces is time-consuming. For instance, the ImageNet have input dimensions of 224×224×3, totaling approximately 0.13 million parameters, and rendering over each dimension iteratively leads to query inefficiency. Standard Simultaneous Perturbation Stochastic Approximation (SPSA) improves efficiency by generating pairs of symmetric vectors and perturbing in multiple directions simultaneously, as follows,
@@
Where $\epsilon$ is a positive scaler and $\xi$ is recommended to follow a symmetric distribution with finite inverse moments (e.g., the Rademacher distribution). The symmetric distribution ensures unbiased exploration of perturbations in both positive and negative directions of parameters at each step. And the finite inverse moments property guarantees that the steps are well-controlled, avoiding excessively large steps due to $\xi^{-1}$ drawn from the distribution (e.g., $\mathbb{E}[1/|\xi|^{p}]$ for some large p), which would otherwise lead to an unstable optimization process. In practical implementations for models with a large number of parameters (e.g., MeZO (Malladi et al., 2023) in large language models (LLMs)), Gaussian noise with zero mean induces substantial perturbations, thereby enhancing exploration across the parameter space and facilitating the escape from local minima. This methodology achieves gradient estimation with only two objective function evaluations, rendering its computational cost independent of input dimensionality. Such computational efficiency has established SPSA as a preferred method for addressing the complexities of high-dimensional deep learning tasks. While increasing $q$ in $q$-SPSA can improve stability in the update direction, setting $q=1$ is sufficient for pretrained LLMs  (Malladi et al., 2023).
2) Descent direction computation. In unconstrained optimization for deep learning, the last gradients $h_{t}$ generally coincide with the estimated ZO gradients $\hat{g_{t}}$ (e.g., ZO-SGD (Ghadimi & Lan, 2013), ZO-SCD (Lian et al., 2016)). To reduce approximation errors, ZO-SGD-sign (Liu et al., 2019) applies an element-wise $sign(\cdot)$ operation. Additionally, ZO-SVRG (Liu et al., 2018), inspired by variance reduction methods in first-order optimization, adjusts the update step by using estimated gradients from previous training examples. CARS (Kim et al., 2021) adaptively selects the smallest function value in each iteration, which helps maintain monotonicity during optimization.
3) Parameter updating. Normally, for most ZO methods, parameters are updated in a similar way with FO optimizers, and the learning rate $\eta_{t}$ is set to constant. Except for the special design for achieving some constraint prerequisites, several methods make an effort to strike a balance between converge speed and accuracy. ZO-AdaMM (Chen et al., 2019) uses an adaptive learning rate and refines gradient estimation by incorporating momentum from past information. This approach is particularly effective in handling complex and evolving optimization landscapes, where the function’s behavior may vary over time or be hard to capture with straightforward gradient approximations.
@@ 3.2 Zeroth-Order Optimization for Catastrophic Forgetting
(a) FO-SGD
(b) ZO-SGD
Figure 3: Trajectory of FO and ZO Optimization during Overcoming Forgetting. The trajectory is taken when using the total loss from both tasks (cyan) and the gradients from each individual task at fixed points during optimization (red and blue). The trends of ZO optimization hold the potential to manage forgetting and learning.
Rationality. ZO optimization leverages the function values of the forward passes to approximate FO gradients, making it feasible to avoid gradient bans. This feature enables seamless integration into common forgetting scenarios, such as CL. We explore it in the following three categories.
@@
where $N_{context}$ represents the number of contexts encountered so far. In Experience Replay (Rolnick et al., 2019b), both components use classification loss based on their respective data distributions, so ZO gradients can be expressed as ^$\hat{\nabla}\mathcal{L}_{cur}$ and ^$\hat{\nabla}\mathcal{L}_{replay}$ respectively. However, in the emerging generative replay workflows (Shin et al., 2017), Equation 2 may introduce additional loss for the training of generators. In this case, the generator can be trained using standard backpropagation or in conjunction with ZO training without FO gradients.
ii) Extension-based methods can be divided into fixed and dynamic architectures. Fixed architectures separate model parameters for specialized context learning, while dynamic architectures expand the model size during adaptation. Both approaches mitigate forgetting from the model’s perspective and enable model-agnostic ZO solutions.
@@
Table 1: ZeroFlow Evaluation on CIFAR-100, ImageNet-A, CUB and OmniBenchmark. This table compares average accuracy, final accuracy, and forgetting measures of 2 models, and 4 forgetting scenarios. More intuitive trend please see Figure 2. Note that all ZO optimization w/ query budget $q=1$, which means that all evaluations use just one forward pass.
@@
As Figure 3, we visualize and compare the optimization trajectories of ZO and FO methods in the learning-memory trade-off dynamics of continual learning. The remarkable similarity between these trajectories motivates our investigation, suggesting ZO optimization’s potential in effectively managing the balance between learning and forgetting.
Potential. The intrinsic optimization mechanism of ZO exhibits particular promise in continual learning scenarios. Intuitively, ZO perturbs parameters using random or coordinate-wise directional vectors and observes changes in the evaluation function, effectively optimizing within a noisy environment. This approach enables small parameter modifications to yield significant impacts on target objectives, resulting in distinctive gradient estimations compared to FO optimization. Notably, while ZO methods do not explicitly incorporate sharpness regularization terms (Foret et al., 2021), they naturally facilitate the exploration of flat regions in parameter space. The influence of optimizing flat regions with ZO approaches in continual learning can be summarized in two main manifolds: (i) For previous tasks, the noise-induced parameter robustness enhances resilience against perturbations from new task adaptation; (ii) For new tasks, empirical evidence suggests that convergence to flat minima generally leads to lower generalization error.
Risk. Although ZO demonstrates superior generalization abilities, its practical performance is limited by optimization strategies and the complexity of the optimization setting. Despite significant efforts to reduce convergence error, optimizing models from scratch in high-dimensional space remains challenging due to slow convergence speed (proportional to the parameter dimension $d$). For instance, origin CGE-based ZO training for a model with 12k parameters takes 70.32 hours in DeepZero (Chen et al., 2024). Such computational demands render from scratch training impractical for high-dimensional CL models, particularly those employing expansion-based architectures. Consequently, we focus our discussion on leveraging ZO optimization to overcome forgetting within a pre-training context.